The ΔG°’ value for the malate DH reaction is +7.10 kcal/mol, and the ΔG°’ value for the citrate synthase reaction is -7.53 kcal/mol. Determine Keq for the combined, coupled reaction (e.g., L-malate <--> citrate, plus ancillary cofactors, reactants, and side-products), and compare this value to Keq for the malate DH reaction alone. By what factor do you have to multiply the Keq of the malate DH reaction to equal the Keq for the coupled reaction. Assume a temperature of 37 °C and report your answer to the nearest ones.
Answer #1

The reaction is: L-malate <------> citrate

With the given ΔG, let's calculate ΔG of the rxn:
ΔGrxn = -7.53 - 7.10 = -14.63 kJ/mol or -14630 J/mol

Now to get K:
Keq = exp(-ΔG/RT)
Keq= exp(14630 / 8.3144 * 37+273)
Keq = 281.82

The Keq for the malate only:
Keq = exp(-7100 / 8.3144 * 310)
Keq = 0.0636

Factor = 281.82 / 0.0636 = 4429

Hope this helps